OpenStudy (anonymous):

30 is 6% of what number ?

OpenStudy (solomonzelman):

\(\large\color{blue}{ N\times 0.6=30}\)

OpenStudy (anonymous):

n*0.06=30

OpenStudy (mathmath333):

it should be \(\large\color{blue}{ N\times 0.06=30}\)

OpenStudy (anonymous):

Thank

sammixboo (sammixboo):

\(\rm \color{#20bd23}{\dfrac{part}{whole}=\dfrac{percent}{100}}\) Part = 30 Whole = Unknown Percent = 6 Plug in the values \(\rm \color{#20bd23}{\dfrac{30}{whole}=\dfrac{6}{100}}\) Now we are finding the whole We can first cross multiply 30 by 100, so what is \(\rm \color{#20bd23}{30\times100}\)?

OpenStudy (anonymous):

of 100 @mycraftcapt

sammixboo (sammixboo):

@mycraftcapt

OpenStudy (anonymous):

30,000

OpenStudy (anonymous):

Right